# Misaki Kojima    Naoki Nishida

### v1

@@ Abstract
An all-path reachability (APR, for short) predicate is a pair of a source set and a target set, which are subsets of an object set. APR predicates have been defined for abstract reduction systems (ARSs, for short) and then extended to logically constrained term rewrite systems (LCTRSs, for short) as pairs of constrained terms that represent sets of terms modeling configurations, states, etc. An APR predicate is said to be partially (or demonically) valid w.r.t. a rewrite system if every finite maximal reduction sequence of the system starting from any element in the source set includes an element in the target set. Partial validity of APR predicates w.r.t. ARSs is defined by means of two inference rules, which can be considered a proof system to construct (possibly infinite) derivation trees for partial validity. On the other hand, a proof system for LCTRSs consists of four inference rules, and thus there is a gap between the inference rules for ARSs and LCTRSs. In this paper, we revisit the framework for APR analysis and adapt it to verification of not only safety but also liveness properties. To this end, we first reformulate an abstract framework for partial validity w.r.t. ARSs so that there is a one-to-one correspondence between the inference rules for partial validity w.r.t. ARSs and LCTRSs. Secondly, we show how to apply APR analysis to safety verification. Thirdly, to apply APR analysis to liveness verification, we introduce a novel stronger validity of APR predicates, called total validity, which requires not only finite but also infinite execution paths to reach target sets. Finally, for a partially valid APR predicate with a cyclic-proof tree, we show a necessary and sufficient condition for the tree to ensure total validity, showing how to apply APR analysis to liveness verification.
@@
abstract reduction system, cyclic proof, constrained rewriting, runtime-error verification
@@ 1 Introduction
Recently, approaches to program verification by means of logically constrained term rewrite systems (LCTRSs, for short) [KN13frocos] are well investigated [FKN17tocl, WM18, CL18, NW18vstte, KN18eptcs, KNS19ss, CLB23, MNKS23wst, KNM25jlamp]. LCTRSs are useful models of not only functional but also imperative programs, which are evaluated sequentially. For instance, equivalence checking by means of LCTRSs is useful to ensure the correctness of terminating functions (cf. [FKN17tocl]). Since the reduction of rewrite systems is in general non-deterministic, rewrite systems are reasonable models of concurrent programs. A transformation of sequential programs into LCTRSs [FKN17tocl, KN18eptcs] has been extended to concurrent programs [KNM25jlamp]. In addition, a method for runtime-error verification by means of all-path reachability analysis of LCTRSs has been developed for, e.g., race and starvation freedom [KN23jlamp, KN23padl, KN24rp].
An all-path reachability (APR, for short) predicate over an object set $A$ is a pair $P\Rightarrow Q$ of a source set $P$ and a target set $Q$, which are subsets of $A$. The APR predicate is said to be partially valid (or demonically valid [CL18]) w.r.t. a rewrite system $\mathcal{S}$ if every finite execution path—a maximal reduction sequence—of $\mathcal{S}$ starting from any element $s$ in $P$ includes an element in $Q$. Partial validity of $P\Rightarrow Q$ w.r.t. $\mathcal{S}$ means that every finite execution from $P$ eventually reaches $Q$. The APR problem w.r.t. $\mathcal{S}$ is a problem to determine whether given APR predicates are partially valid w.r.t. $\mathcal{S}$ or not.
@@
$\mbox{\rm{\rm{Subsumption}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $P\subseteq Q$}\qquad\mbox{\rm{\rm{Step}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \partial_{\mathcal{A}}(P\setminus Q)\Rightarrow Q\penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $(P\setminus Q)$ is $\mathcal{A}$-runnable }$
Figure 1: Inference rules of $\mathsf{DVP}$ [CL18] for partial validity of APR predicates w.r.t. an an ARS $\mathcal{A}=(A,\to_{\mathcal{A}})$, where $\partial_{\mathcal{A}}(P)=\{t^{\prime}\in A\mid\exists t\in P.\ t\to_{\mathcal{A}}t^{\prime}\}$.
The abstract APR framework has been adapted to LCTRSs, and the proof system $\mathsf{DCC}$ for partial validity has been introduced [CL18]. An APR predicate over a signature $\Sigma$ is a pair $\langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle$ of constrained terms $\langle s\,|\,\phi\rangle,\langle t\,|\,\psi\rangle$ over $\Sigma$, which represents sets of configurations, states, etc. To simplify the discussion, we assume that the two constrained terms have no shared variables. Note that $\mathcal{R}$ is an ARS for the set of terms over $\Sigma$, and a constrained term $\langle s\,|\,\phi\rangle$ can be considered sets of ground normalized instances of $s$ by means of substitutions satisfying $\phi$. The proof system $\mathsf{DCC}$ w.r.t. an LCTRS $\mathcal{R}$ consists of the four inference rules shown in Figure˜2. Given a set $G$ of APR predicates over $\Sigma$, if we succeed in constructing proof trees for all APR predicates in $G$ by $\mathsf{DCC}$ w.r.t. $\mathcal{R}$, then all APR predicates in $G$ are partially valid w.r.t. $\mathcal{R}$. Note that the application of Subs is followed by Circ or either Axiom or Der.
@@
$\mbox{\rm{\rm{Der}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \langle s_{1}\,|\,\phi_{1}\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ \,\ldots\penalty 10000\ \,\langle s_{n}\,|\,\phi_{n}\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }{\penalty 10000\ \langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }\penalty 10000\ \mbox{if $\langle s\,|\,\phi\rangle$ is $\mathcal{R}$-runnable, }$
where $\Delta_{\mathcal{R}}(\langle s\,|\,\phi\rangle)=\{\langle s_{i}\,|\,\phi_{i}\rangle\mid 1\leq i\leq n\}$ for some $n>0$, and $\Delta_{\mathcal{R}}(\langle s^{\prime}\,|\,\phi^{\prime}\rangle):=\{\langle s^{\prime\prime}\,|\,\phi^{\prime\prime}\rangle\mid\langle s^{\prime}\,|\,\phi^{\prime}\rangle\mathrel{\leadsto_{\mathcal{R}}}\langle s^{\prime\prime}\,|\,\phi^{\prime\prime}\rangle\}$.
$\mbox{\rm{\rm{Circ}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \langle t^{\prime}\,|\,\phi\wedge\phi^{\prime\prime}\wedge\psi^{\prime}\rangle\Rightarrow\langle t\,|\,\psi\rangle\quad\langle s\,|\,\phi\wedge\neg\phi^{\prime\prime}\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }{\penalty 10000\ \langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }$
where $\langle s^{\prime}\,|\,\phi^{\prime}\rangle\Rightarrow\langle t^{\prime}\,|\,\psi^{\prime}\rangle$ is a freshly renamed APR predicate in $G$, $\{\vec{x}\}=\mathcal{V}ar(s^{\prime},\phi^{\prime})$, and $\phi^{\prime\prime}=(\exists\vec{x}.\ (s=s^{\prime})\wedge\phi^{\prime})$.
@@
The two proof systems $\mathsf{DVP}$ and $\mathsf{DCC}$ have no one-to-one correspondence between their inference rules. The main role of Circ in $\mathsf{DCC}$ is the introduction of circularity to proof trees of $\mathsf{DCC}$ so as to make the trees finite as in cyclic proofs [Bro05]. Since the initial role of $\mathsf{DVP}$ is to define partial validity, $\mathsf{DVP}$ does not consider circularity and is used to construct possibly infinite proof trees as in $\mbox{LKID}^{\omega}$ [Bro06phd, Chapter 4]. When we do not use Circ in $\mathsf{DCC}$ and allow to construct infinite proof trees, roughly speaking, rules Axiom, Subs, and Der in $\mathsf{DCC}$ correspond to rules Subsumption and Step in $\mathsf{DVP}$ as follows:
  • The application of Axiom and Der in $\mathsf{DCC}$ is simulated by Subsumption and Step in $\mathsf{DVP}$, respectively;
  • the sequential application of Subs and Axiom in $\mathsf{DCC}$ is simulated by Subsumption in $\mathsf{DVP}$, and vice versa;
  • the sequential application of Subs and Der in $\mathsf{DCC}$ is simulated by Step in $\mathsf{DVP}$, and vice versa.
This gap between the proof systems—the absence of one-to-one correspondence between the inference rules—makes the correctness proof of $\mathsf{DCC}$ non-trivial. If $\mathsf{DCC}$ is clearly an instance of $\mathsf{DVP}$, then the correctness of $\mathsf{DCC}$ is immediate from that of $\mathsf{DVP}$; it suffices to show that rules in $\mathsf{DCC}$ are instances of corresponding rules in $\mathsf{DVP}$. In addition, the abstract framework is sufficiently useful to investigate APR-based approaches to verification of, e.g., safety and liveness properties because concurrent transition systems which can be represented by ARSs are usual models for such properties.
@@
The APR framework seems to be well suited to verification of liveness properties: We let $Q$ be a set of states to be eventually reached; we let $P_{0}$ be either the set of initial states or a set of intermediate states to reach states in $Q$; then, we attempt to prove partial validity of $P_{0}\Rightarrow Q$; the difference operation $P\setminus Q$ remains states in $P$ that have not yet reached any state in $Q$. A weakness of partial validity of APR predicates is that infinite execution paths do not have to be considered, while it does not matter for verification of safety properties: Roughly speaking, reduction sequences ending with error states of safety properties are finite execution paths, and infinite execution paths can be excluded. However, for general runtime-error verification, not only terminating executions but also non-terminating ones have to be taken into account because a runtime error to be verified may happen in a non-terminating execution. For this problem, in [KN23jlamp], simpler APR predicates have been introduced as mentioned above, and a given LCTRS is modified in order to make all finite prefixes of (possibly infinite) execution paths finite execution paths. Unfortunately, the limitation of APR predicates eliminates the good compatibility that exists between $P\setminus Q$ and liveness verification.
In this paper, we aim to develop an abstract foundation for APR analysis toward runtime-error verification. To this end, we revisit the framework for partial validity of APR analysis and adapt it to verification of not only safety but also liveness properties such as race freedom and starvation freedom. We first reformulate an abstract framework for partial validity of APR predicates w.r.t. ARSs so that there is a one-to-one correspondence between the inference rules for partial validity w.r.t. ARSs and LCTRSs. To be more precise, we reformulate inference rules for partial validity of APR predicates w.r.t. ARSs (Section˜3), and adapt a cyclic-proof system for partial validity w.r.t. LCTRSs to ARSs (Section˜4). Secondly, we show how to apply APR analysis to safety verification (Section˜5). Thirdly, to apply APR analysis to liveness verification, we introduce a novel stronger validity of APR problems, called total validity, which requires not only finite but also infinite execution paths to reach target sets (Section˜6.1). Finally, for a partially valid APR predicate with a cyclic-proof tree, we show a sufficient condition for the tree to ensure total validity, showing how to apply APR analysis to liveness verification (Section˜6.2). All omitted proofs of the claims are provided in both the appendix.
@@
Since “demonical validity” [CL18] (partial validity in this paper) does not consider infinite execution paths and we consider a stronger validity for all execution paths, we renamed “demonical validity” to “partial validity”, which originates from partial and total correctness of programs. Notice that total correctness of programs requires termination of programs, but total validity does not require termination but requires all (possibly infinite) execution paths reach elements in target sets.
@@ 2 Preliminaries
The derivative of a set $P$ ($\subseteq A$) is the set $\partial_{\mathcal{A}}(P)=\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}t\}$. The set $P$ is called runnable w.r.t. $\mathcal{A}$ ($\mathcal{A}$-runnable, for short) if $P\neq\emptyset$ and $P\cap\mathit{NF\!}_{\mathcal{A}}=\emptyset$. By definition, it is clear that $\partial_{\mathcal{A}}(P)\neq\emptyset$ if and only if $P\setminus\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$.
We define execution paths of ARSs.
@@
Note that execution paths are defined coinductively by using two rules in $\mathsf{DVP}$ w.r.t. $\mathcal{A}$ [CL18] (Figure˜1 in this paper).
@@
An APR predicate $P\Rightarrow Q$ over $A$ is said to be partially valid (demonically valid [CL18]) w.r.t. $\mathcal{A}$ if each execution path starting from an element in $P$ either eventually reaches an element in $Q$ or is infinite.
Partial validity is defined coinductively as follows.
@@
An APR predicate $P\Rightarrow Q$ over $\mathcal{A}$ is said to be partially valid w.r.t. $\mathcal{A}$, written as $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$, if $(P\Rightarrow Q)\in\nu\widehat{\mathsf{DVP}(\mathcal{A})}$, where $\mathsf{DVP}(\mathcal{A})$ consists of the rules Subsumption and Step in Figure˜1. We write $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ if $P\Rightarrow Q$ is not partially valid w.r.t. $\mathcal{A}$.
Regarding partial validity, APR predicates have the following properties.
@@
Proof 2.5.
Trivial by definition.
For an APR predicate $P\Rightarrow Q$, regarding partial validity, Proposition˜2.4 (1 and 2) implies soundness and completeness of splitting $P$ to $n$ sets $P_{1},\ldots,P_{n}$ (i.e., $P=P_{1}\cup\cdots\cup P_{n}$):
  • $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q_{i}$ for all $1\leq i\leq n$ if and only if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}(P_{1}\cup\cdots\cup P_{n})\Rightarrow Q$ (Proposition˜2.4-12).
On the other hand, splitting the target set $Q$ to $Q_{1},\ldots,Q_{n}$ (i.e., $Q=Q_{1}\cup\cdots\cup Q_{n}$) is sound but not complete in general:
  • if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q_{i}$ for all $1\leq i\leq n$, then, $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow(Q_{1}\cup\cdots\cup Q_{n})$ (Proposition˜2.4-1), but
  • the other direction does not hold in general.
@@
3 Reformulation of Inference Rules fo ARSs
In this section, we reformulate the proof system $\mathsf{DVP}$ for partial validity w.r.t. ARSs in [CL18] to establish a one-to-one correspondence between the proof rules for partial validity w.r.t. ARSs and LCTRSs.
By abstracting the inference rules in Figure˜2, we obtain inference rules for ARSs, which have one-to-one correspondence with the rules for LCTRSs.

### v2

@@ Abstract
An all-path reachability (APR, for short) predicate over an object set is a pair of a source set and a target set, which are subsets of the object set. APR predicates have been defined for abstract reduction systems (ARSs, for short) and then extended to logically constrained term rewrite systems (LCTRSs, for short) as pairs of constrained terms that represent sets of terms modeling configurations, states, etc. An APR predicate is said to be partially (or demonically) valid w.r.t. a rewrite system if every finite maximal reduction sequence of the system starting from any element in the source set includes an element in the target set. Partial validity of APR predicates w.r.t. ARSs is defined by means of two inference rules, which can be considered a proof system to construct (possibly infinite) derivation trees for partial validity. On the other hand, a proof system for LCTRSs consists of four inference rules, leaving a gap between the inference rules for ARSs and LCTRSs. In this paper, we revisit the framework for APR analysis and adapt it to verification of not only safety but also liveness properties. To this end, we first reformulate an abstract framework for partial validity w.r.t. ARSs so that there is a one-to-one correspondence between the inference rules for partial validity w.r.t. ARSs and LCTRSs. Secondly, we show how to apply APR analysis to safety verification. Thirdly, to apply APR analysis to liveness verification, we introduce a novel stronger validity of APR predicates, called total validity, which requires not only finite but also infinite execution paths to reach target sets. Finally, for a partially valid APR predicate with a cyclic-proof tree, we show a necessary and sufficient condition for the tree to ensure total validity. The condition implies that if there exists a cyclic-proof tree for an APR predicate, the proof graph of which is acyclic, then the APR predicate is totally valid.
@@
abstract reduction system, reachability, cyclic proof, runtime-error verification
@@ 1 Introduction
Recently, approaches to program verification by means of logically constrained term rewrite systems (LCTRSs, for short) [KN13frocos] are well investigated [FKN17tocl, WM18, CL18, NW18vstte, KN18eptcs, KNS19ss, CLB23, MNKS23wst, KNM25jlamp]. LCTRSs are useful models of not only functional but also imperative programs. For instance, equivalence checking by means of LCTRSs is useful to ensure the correctness of terminating functions (cf. [FKN17tocl]). Since the reduction of rewrite systems is in general non-deterministic, rewrite systems are reasonable models of concurrent programs. A transformation of sequential programs into LCTRSs [FKN17tocl, KN18eptcs] has been extended to concurrent programs [KNM25jlamp]. In addition, a method for runtime-error verification by means of all-path reachability analysis of LCTRSs has been developed for, e.g., race and starvation freedom [KN23jlamp, KN23padl, KN24rp].
An all-path reachability (APR, for short) predicate over an object set $A$ is a pair $P\Rightarrow Q$ of a source set $P$ and a target set $Q$, which are subsets of $A$. The APR predicate is said to be partially valid (or demonically valid [CL18]) w.r.t. a rewrite system $\mathcal{A}$ if every finite execution path—a maximal reduction sequence—of $\mathcal{A}$ starting from any element $s$ in $P$ includes an element in $Q$. Partial validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ means that every finite execution from $P$ eventually reaches $Q$. The APR problem w.r.t. $\mathcal{A}$ is a problem to determine whether given APR predicates are partially valid w.r.t. $\mathcal{A}$ or not.
@@
$\mbox{\rm{\rm{Subsumption}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $P\subseteq Q$.}\qquad\mbox{\rm{\rm{Step}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \partial_{\mathcal{A}}(P\setminus Q)\Rightarrow Q\penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $(P\setminus Q)$ is $\mathcal{A}$-runnable. }$
Figure 1: Inference rules of $\mathsf{DVP}$ [CL18] for partial validity of APR predicates w.r.t. an ARS $\mathcal{A}=(A,\to_{\mathcal{A}})$, where $\partial_{\mathcal{A}}(P)=\{t^{\prime}\in A\mid\exists t\in P.\ t\to_{\mathcal{A}}t^{\prime}\}$.
The abstract APR framework has been adapted to LCTRSs, and the proof system $\mathsf{DCC}$ for partial validity has been presented [CL18]. An APR predicate over a signature $\Sigma$ is a pair $\langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle$ of constrained terms $\langle s\,|\,\phi\rangle,\langle t\,|\,\psi\rangle$ over $\Sigma$, which represents sets of configurations, states, etc. To simplify the discussion, we assume that the two constrained terms have no shared variables. Note that $\mathcal{R}$ is an ARS for the set of terms over $\Sigma$, and a constrained term $\langle s\,|\,\phi\rangle$ can be considered sets of ground normalized instances of $s$ by means of substitutions satisfying $\phi$. The proof system $\mathsf{DCC}$ w.r.t. an LCTRS $\mathcal{R}$ consists of the four inference rules shown in Figure˜2. Given a set $G$ of APR predicates over $\Sigma$, if we succeed in constructing proof trees for all APR predicates in $G$ by $\mathsf{DCC}$ w.r.t. $\mathcal{R}$, then all APR predicates in $G$ are partially valid w.r.t. $\mathcal{R}$. Note that the application of Subs is followed by Circ or either Axiom or Der.
@@
$\mbox{\rm{\rm{Der}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \langle s_{1}\,|\,\phi_{1}\rangle\Rightarrow\langle t\,|\,\psi\rangle\quad\ldots\quad\langle s_{n}\,|\,\phi_{n}\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }{\penalty 10000\ \langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }\penalty 10000\ \mbox{if $\langle s\,|\,\phi\rangle$ is $\mathcal{R}$-runnable, }$
where $\{\langle s_{i}\,|\,\phi_{i}\rangle\mid 1\leq i\leq n\}=\{\langle s^{\prime}\,|\,\phi^{\prime}\rangle\mid\langle s\,|\,\phi\rangle\mathrel{\leadsto_{\mathcal{R}}}\langle s^{\prime}\,|\,\phi^{\prime}\rangle\}$ for some $n>0$.
$\mbox{\rm{\rm{Circ}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \langle t^{\prime}\,|\,\phi\wedge\phi^{\prime\prime}\wedge\psi^{\prime}\rangle\Rightarrow\langle t\,|\,\psi\rangle\qquad\langle s\,|\,\phi\wedge\neg\phi^{\prime\prime}\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }{\penalty 10000\ \langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }$
where $(\langle s^{\prime}\,|\,\phi^{\prime}\rangle\Rightarrow\langle t^{\prime}\,|\,\psi^{\prime}\rangle)\in G$, $\{\vec{x}\}=\mathcal{V}ar(s^{\prime},\phi^{\prime})$, and $\phi^{\prime\prime}=(\exists\vec{x}.\ (s=s^{\prime})\wedge\phi^{\prime})$.
@@
The two proof systems $\mathsf{DVP}$ and $\mathsf{DCC}$ have no one-to-one correspondence between their inference rules. The main role of Circ in $\mathsf{DCC}$ is the introduction of circularity to proof trees of $\mathsf{DCC}$ so as to make the trees finite as in cyclic proofs [Bro05]. Since the initial role of $\mathsf{DVP}$ is to define partial validity, $\mathsf{DVP}$ does not consider circularity and is used to construct possibly infinite proof trees as in $\mbox{LKID}^{\omega}$ [Bro06phd, Chapter 4]. When we do not use Circ in $\mathsf{DCC}$ and allow to construct infinite proof trees, roughly speaking, rules Axiom, Subs, and Der in $\mathsf{DCC}$ correspond to rules Subsumption and Step in $\mathsf{DVP}$, while there is no one-to-one correspondence between the inference rules in $\mathsf{DCC}$ and $\mathsf{DVP}$. The absence of one-to-one correspondence makes the correctness proof of $\mathsf{DCC}$ non-trivial. If $\mathsf{DCC}$ is clearly an instance of $\mathsf{DVP}$, then the correctness of $\mathsf{DCC}$ is immediate from that of $\mathsf{DVP}$; it suffices to show that rules in $\mathsf{DCC}$ are instances of corresponding rules in $\mathsf{DVP}$. In addition, the abstract framework is sufficiently useful to investigate APR-based approaches to verification of, e.g., safety and liveness properties, because concurrent transition systems which can be represented by ARSs are usual models for such properties.
In this paper, we aim to develop an abstract foundation for APR analysis toward runtime-error verification. To this end, we revisit the framework for partial validity of APR analysis and adapt it to verification of not only safety but also liveness properties such as race and starvation freedom.
@@
The APR framework seems to be well suited to verification of liveness properties: We let $P_{0}$ be either the set of initial states or a set of intermediate states to reach states in a target set $Q$; then, we attempt to prove partial validity of $P_{0}\Rightarrow Q$; the difference operation $P\setminus Q$ remains states in $P$ that have not yet reached any state in $Q$. The shortcoming of partial validity is that infinite execution paths do not have to be considered, while it does not matter for verification of safety properties: Roughly speaking, reduction sequences ending with error states of safety properties are finite execution paths, and infinite execution paths can be excluded. However, for general runtime-error verification, not only terminating executions but also non-terminating ones have to be taken into account because a runtime error to be verified may happen in a non-terminating execution. For this problem, in [KN23jlamp], simpler APR predicates have been introduced as mentioned above, and a given LCTRS is modified in order to make all finite prefixes of (possibly infinite) execution paths finite execution paths of the modified LCTRS. Unfortunately, the limitation of APR predicates eliminates the good compatibility that exists between $P\setminus Q$ and liveness verification.
We first reformulate an abstract framework for partial validity of APR predicates w.r.t. ARSs so that there is a one-to-one correspondence between the inference rules for partial validity w.r.t. ARSs and LCTRSs. To be more precise, we reformulate inference rules for partial validity w.r.t. ARSs (Section˜3), and adapt a cyclic-proof system for partial validity w.r.t. LCTRSs in [KN24jip, Section 3.2] to ARSs (Section˜4). Secondly, we show how to apply APR analysis to safety verification (Section˜5). Thirdly, to apply APR analysis to liveness verification, we introduce a novel stronger validity of APR problems, called total validity, which requires not only finite but also infinite execution paths to reach target sets (Section˜6.1). Finally, for a partially valid APR predicate with a cyclic-proof tree, we show a necessary and sufficient condition for the tree to ensure total validity, showing how to apply APR analysis to liveness verification (Section˜6.2). All omitted proofs of the claims are provided in the appendix.
@@
Since “demonical validity” [CL18] (partial validity in this paper) does not consider infinite execution paths and we consider a stronger validity for all execution paths, we renamed “demonical validity” to “partial validity”, which originates from partial and total correctness of programs. Notice that total correctness of programs requires termination of programs, but total validity does not require termination but requires all (possibly infinite) execution paths to reach elements in target sets.
@@ 2 Preliminaries
The derivative of a set $P$ ($\subseteq A$) is the set $\partial_{\mathcal{A}}(P)=\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}t\}$. The set $P$ is called runnable w.r.t. $\mathcal{A}$ ($\mathcal{A}$-runnable, for short) if $P\neq\emptyset$ and $P\cap\mathit{NF\!}_{\mathcal{A}}=\emptyset$. By definition, it is clear that $\partial_{\mathcal{A}}(P)\neq\emptyset$ if and only if $P\setminus\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$. We define execution paths of ARSs.
@@
Note that execution paths are defined coinductively by using two inference rules in $\mathsf{DVP}$ w.r.t. $\mathcal{A}$ [CL18] (Figure˜1 in this paper).
@@
An APR predicate $P\Rightarrow Q$ over $A$ is said to be partially valid (or demonically valid [CL18]) w.r.t. $\mathcal{A}$ if each execution path starting from an element in $P$ either eventually reaches an element in $Q$ or is infinite. Partial validity is defined coinductively as follows.
@@
An APR predicate $P\Rightarrow Q$ over $A$ is said to be partially valid w.r.t. $\mathcal{A}$, written as $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$, if $(P\Rightarrow Q)\in\nu\widehat{\mathsf{DVP}(\mathcal{A})}$, where $\mathsf{DVP}(\mathcal{A})$ consists of the rules Subsumption and Step in Figure˜1. We write $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ if $P\Rightarrow Q$ is not partially valid w.r.t. $\mathcal{A}$.
By definition, APR predicates trivially have the following properties for partial validity.
@@
For an APR predicate $P\Rightarrow Q$, regarding partial validity, Proposition˜2.4 12 implies soundness and completeness of splitting $P$ to $n$ sets $P_{1},\ldots,P_{n}$ (i.e., $P=P_{1}\cup\cdots\cup P_{n}$): $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q_{i}$ for all $1\leq i\leq n$ if and only if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}(P_{1}\cup\cdots\cup P_{n})\Rightarrow Q$. On the other hand, splitting the target set $Q$ to $Q_{1},\ldots,Q_{n}$ (i.e., $Q=Q_{1}\cup\cdots\cup Q_{n}$) is sound by Proposition˜2.4 1, but not complete in general: if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q_{i}$ for all $1\leq i\leq n$, then, $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow(Q_{1}\cup\cdots\cup Q_{n})$, but the other direction does not hold in general.
@@
3 Reformulation of Inference Rules for ARSs
In this section, we reformulate the proof system $\mathsf{DVP}$ for partial validity w.r.t. ARSs to establish a one-to-one correspondence between the inference rules for ARSs and LCTRSs.
By abstracting the inference rules for LCTRSs in Figure˜2, we obtain inference rules for ARSs, which have one-to-one correspondence with the inference rules for LCTRSs.